### v1

# Ceresa cycles of $X_{0}(N)$

Abstract.
The Ceresa cycle is an algebraic 1-cycle on the Jacobian of an algebraic curve. Although it is homologically trivial, Ceresa famously proved that for a very general complex curve of genus at least 3, it is non-trivial in the Chow group. In this paper we study the Ceresa cycle attached to the complete modular curve $X_{0}(N)$. For prime level $p$ we give a complete description, namely we prove that if $X_{0}(p)$ is not hyperelliptic, then its Ceresa cycle is non-torsion. For general level $N$, we prove that there are finitely many $X_{0}(N)$ with torsion Ceresa cycle. Our method relies on the relationship between the vanishing of the Ceresa cycle and Chow-Heegner points on the Jacobian. We use the geometry and arithmetic of modular Jacobians to prove that such points are of infinite order and therefore deduce non-vanishing of the Ceresa cycle.
Key words and phrases:
Modular Curves, Ceresa Cycle, Gross-Kudla-Schoen modified diagonal cycle, Chow-Heegner Divisor
2000 Mathematics Subject Classification:
14C25, 14G35, 11G18
1. Introduction
Let $C$ be a smooth, projective and geometrically integral curve, defined over an algebraically closed field $K$. Fix a degree $1$ base point divisor $e$ and let $\iota_{e}$ be the corresponding Abel-Jacobi map, which embeds the curve into its Jacobian $J$. The Ceresa cycle associated to $(C,e)$ is classically defined as
$\operatorname{Cer}(C,e):=[\iota_{e}(C)]-[-1]^{*}[\iota_{e}(C)]$
which is an element of the Chow group $CH_{1}(J)$. This cycle is always homologically trivial, that is, it lies in the kernel of the cycle class map form $CH_{1}(J)$ into a Weil cohomology theory, but crucially it is not always trivial in the Chow group. In fact, for a very general complex curve of genus at least 3, Ceresa [7] proved that it is algebraically, and in particular rationally, non-trivial independent of the choice of base point divisor.
In the case of hyperelliptic curves, taking $e$ to be a Weierstrass point shows that $\operatorname{Cer}(C,e)=0$. Few other explicit examples are known. Some families with torsion Ceresa cycle have been studied by Qiu and Zhang [35], [34]. Buhler, Schoen and Top [6] studied the Ceresa cycle in conjunction with the Beilinson-Bloch conjectures. Laga and Shnidman [25] treated the case of bielliptic Picard curves. In recent work, Ellenberg, Logan and Srinivasan [16] developed an algorithm for verifying non-vanishing of the Ceresa cycle, which is guaranteed to terminate if the Sato-Tate group is GSp. The Ceresa cycle has been widely studied for Fermat curves and their quotients [24], [36], [32], [27], [4], [21]. In particular, Eskandari and Murty [17] proved that for any prime $p>7$, the Ceresa cycle associated to the Fermat curve $x^{p}+y^{p}=z^{p}$ is non-torsion for any choice of base point. The proof relies on the close relation between Chow-Heegner divisors, first constructed by Darmon, Rotger and Sols [10], and vanishing of the Ceresa cycle; a relation which is also crucial in our argument.
In this paper we study a classically interesting family of curves, namely the modular curves whose non-cuspidal points parametrise isomorphism classes of elliptic curves with a marked isogeny. We study the Ceresa cycle of $X_{0}(N)$ inside $\operatorname{CH}_{1}(J)\otimes\mathbb{Q}$, henceforth simply $\operatorname{CH}_{1}(J)$. As detailed in Theorem 3 (c.f. [39]), the Ceresa cycle can only possibly vanish for one choice of basepoint divisor class, namely $e=\frac{K_{C}}{2g-2}$, where $K_{C}$ is the canonical class and $g$ is the genus of the curve. For the rest of this paper we write $\operatorname{Cer}(N)$ for $\operatorname{Cer}\left(X_{0}(N),\frac{K_{C}}{2g-2}\right)$. Recently, a large family of modular curves was studied by Kerr, Li, Qiu and Yang [23]. This work proves that the Ceresa cycle associated to the complete modular curve corresponding to the congruence subgroup $\Gamma(2)\cap\Gamma_{1}(2N)$ is non-zero in $\operatorname{CH}_{1}(J)$, for $N$ sufficiently large. We remark that their results are implied by our work, combined with a statement about Ceresa cycles and covers (see Proposition 24). The main results of our paper are the following.
Theorem 1.
The Ceresa cycle $\operatorname{Cer}(p)$ is non-zero if and only if $X_{0}(p)$ is not hyperelliptic. That is, if $p>71$ or $p\in\{43,53,61,67\}$, then $\operatorname{Cer}(p)$ is non-trivial.
Our second result deals with the case of arbitrary level $N$ and has significantly more exclusions.
Theorem 2.
There are finitely many $N$ such that $\operatorname{Cer}(N)=0$. More precisely, if
$N>2^{5}\times 3^{4}\times 5^{2}\times 7^{2}\times\displaystyle\prod_{\begin{subarray}{c}11\leq p\leq 71,\\ p\not\in\{43,53,61,67\}\end{subarray}}p$
then $\operatorname{Cer}(N)$ is non-zero in $CH_{1}(J)$.
We note that Theorem 2 combined with Theorem 3 (reproduced from [39]) proves a significant part of [33, Conjecture 1.2.2].
As alluded to previously, the key ingredient of our proof is the relation between the Ceresa cycle on $X_{0}(N)$ and certain (scaled) Chow-Heegner divisors associated to Hecke operators. More specifically, we are able to reduce the non-vanishing of the Ceresa cycle to proving that a carefully constructed rational point on the Jacobian is non-torsion. The study of rational points on Jacobian of modular curves has been of great arithmetic interest. The theorems of Manin [28] and Drinfeld [15] show that the subgroup of the Jacobian generated by linear equivalences of differences of cusps is finite. Mazur [29] showed that for prime level $p\geq 5$ the rational torsion subgroup of $J_{0}(p)$ is generated by the differences of the two cusps of $X_{0}(p)$. In fact, this is conjectured to hold more generally for composite $N$. Moreover, we also have some understanding of non-torsion points of $J_{0}(N)(\mathbb{Q})$, in large part due Gross and Zagier’s classical study of Heegner divisors [20].
Our methodology is similar to [23], as we reduce non-vanishing to the study of a certain divisor in the modular Jacobian. However, the key difference is that the congruence subgroup considered in the previous work allows the authors to reduce to Heegner divisors and conclude that such points are of infinite order by using the celebrated Gross-Zagier formula to show that the Néron-Tate height of the divisor is non-zero. In our case, the associated points do not fall into a pre-existing framework, and we instead make use of the arithmetic and geometry of the modular Jacobian to deduce that the associated point is non-torsion. We do not believe such points have received in-depth study in the modular curves literature to date, as despite being linear combinations of Heegner divisors, they do not satisfy the assumptions of the Gross-Zagier height theorem [20].
One drawback of our method (and of [23]) is that it does not apply to examples where $J_{0}(N)(\mathbb{Q})$ has rank 00. A famous examples of this is $X_{0}(64)$, which is simply the Fermat quartic, whose Jacobian has $\mathbb{Q}$-rank zero, but for which the Ceresa cycle does not vanish [21]. There are other examples excluded by bound of Theorem 2 with finite $J_{0}(N)(\mathbb{Q})$, for instance $N=34,38,51,55\ldots$. The curve $X_{0}(34)$ has genus $3$ and we were able to verify non-vanishing using [16], however we were unable to apply this to higher genus curves due to the automorphism hypothesis. Moreover, our examples fail the cohomological vanishing criteria proved in [26].
While it is seemingly possible to use the methods of Section 5 to treat all levels $N$ for which $X_{0}(N)$ is non-hyperelliptic and $J_{0}(N)(\mathbb{Q})$ has positive rank, there are sufficiently many cases to consider that it is impractical and too computationally expensive at this time.
Beyond constructing new points of infinite order on the modular Jacobian, the main arithmetic application of our work is towards triple product $L$-functions. Namely, there are relations between the height of $(f,g,h)$-isotypic Hecke components of the Gross-Kudla-Schoen cycle (an object closely related to the Ceresa cycle e.g. [39]), and the special value $L^{\prime}(f\times g\times h,2)$, by analogy with the Gross-Zagier height formula [20] [38] [2]. Moreover, similar points are also studied by Dogra and Le Fourn [14] in the context of the quadratic Chabauty method for modular curves.
This paper is organised as follows. In Section 2, we recall some standard constructions associated to Chow groups and define ‘shadow points’ corresponding to Hecke operators on $J_{0}(N)$. In Section 3, we give a brief overview of some well-known results on the geometry and arithmetic of modular curves which allow us to study the shadow points. In Sections 4 and 5, we prove Theorems 1 and 2 respectively.
Some cases in our analysis required the use of computer calculation. The source code for these is available at https://github.com/jameswrawson/ModularCeresa.
Acknowledgements.
The authors thank Vladimir Dokchitser, Jef Laga, Samir Siksek and Jan Vonk for helpful conversations. Both authors are grateful for the financial support provided by the UK Engineering and Physical Sciences Research Council. The first named author is supported by the EPSRC Doctoral Prize fellowship EP/W524335/1. The second named author is supported by the EPSRC studentship EP/W523793/1.
2. Gross-Kudla-Schoen Cycles and Chow-Heegner Points
In what follows $C$ and $D$ are smooth, projective and geometrically integral curves, defined over an algebraically closed field $K$. We write $J_{C}$ for the Jacobian variety of $C$, and we denote the genus of $C$ by $g$.
2.1. The Ceresa Cycle and Gross-Kudla-Schoen Classes
Another cycle that can be written down for an arbitary curve $C$ and degree 1 divisor $e$ on $C$ is the following.
Definition 1.
The Gross-Kudla-Schoen (modified) diagonal cycle, $\Gamma^{3}(C,e)\in\operatorname{CH}_{1}(C^{3})$, is defined to be
$\Delta^{3}_{C}-\operatorname{pr}_{1,2}^{*}(\Delta_{C})\cdot\operatorname{pr}^{*}_{3}e-\operatorname{pr}^{*}_{1,3}(\Delta_{C})\cdot\operatorname{pr}_{2}^{*}e-\operatorname{pr}^{*}_{2,3}(\Delta_{C})\cdot\operatorname{pr}_{1}^{*}e+\operatorname{pr}_{1,2}^{*}(e\times e)+\operatorname{pr}_{1,3}^{*}(e\times e)+\operatorname{pr}_{2,3}^{*}(e\times e),$
where $\operatorname{pr}_{i}$ denotes the projection on to the $i$th factor, $\operatorname{pr}_{i,j}$ is the projection onto the product of the $i$th and $j$th factor, $\Delta_{C}^{3}$ denotes the small diagonal in $C^{3}$ (i.e. the locus where all 3 points are equal), and $\Delta_{C}$ the diagonal in $C^{2}$.
The following theorem of Zhang [39] relates the vanishing of the Ceresa cycle to the vanishing of the Gross-Kudla-Schoen diagonal cycle.
Theorem 3 (Zhang).
If $\Gamma^{3}(C,e)=0$, then $(2g-2)e=K_{C}$, and $\operatorname{Cer}(C,e)=0$. Conversely, if $\operatorname{Cer}(C,e)=0$, then $\Gamma^{3}(C,e)=0$.
In light of this result, we restrict to the case $e=\frac{K_{C}}{2g-2}$ and write $\operatorname{Cer}(C)$ for $\operatorname{Cer}(C,\frac{K_{C}}{2g-2})$, for any choice of canonical divisor.
2.2. Chow-Heegner and Shadow Points
Recall that to any morphism of Jacobians, $\phi:J_{C}\to J_{D}$, there is a corresponding correspondence, $C\leftarrow X\rightarrow D$, which realises the isogeny via push-forward and pullback. When $\phi$ is an endomorphism, this correspondence can be naturally embedded in $C^{2}$, and we may refer to fixed points of $\phi$ as the intersection of the correspondence with the diagonal.
Definition 2.
Let $C$ be a curve of genus $g$, $J_{C}$ its Jacobian and $\phi$ an endomorphism of $J_{C}$. The shadow of $\phi$ is the degree 00 divisor on $C$
$\operatorname{Sh}(\phi)=(2g-2)F_{\phi}-\deg(F_{\phi})K_{C}-\phi(K_{C})-\phi^{\vee}(K_{C})+(\deg(\phi)+\deg(\phi^{\vee}))K_{C}$
where $F_{\phi}$ is the fixed point divisor of $\phi$, $K_{C}$ is a canonical divisor on $C$ and $\phi^{\vee}$ denotes the correspondence traversed in the opposite direction.
We remark that this divisor is a scaling of the corresponding Chow-Heegner point, see [10], [9] for details on the construction of such points. The terminology of “shadows” is adopted from [16], both due to the similarity of the construction, and to avoid any notational ambiguity with the Chow group. If $\phi$ is defined over $\mathbb{Q}$, then so is $\operatorname{Sh}(\phi)$.
The following proposition is similar to those found in [16] and [23].
Proposition 4.
Suppose $J_{C}$ has a non-trivial endomorphism,$\phi$. If the Ceresa cycle is zero in $\operatorname{CH}_{1}(J)$, then $[\operatorname{Sh}(\phi)]\in J_{C}$ is torsion.
Proof.
As mentioned before this proposition, associated to $\phi$ there is a natural class $X_{\phi}$ in $C^{2}$, given by the image of the correspondence realising $\phi$. We can push this class forward through $\Gamma^{3}(C,e)$, by the following $\operatorname{pr}_{3,*}\left(\Gamma^{3}(C,e)\cdot\operatorname{pr}_{1,2}^{*}X_{\phi}\right)$, where $e=\frac{K_{C}}{2g-2}$. This can be calculated by computing this pushforward for each term in the definition of $\Gamma^{3}(C,e)$. We summarise these calculations in Table 1, using the notation of the previous definition.
$z\in CH_{1}(C^{3})$ $z\cdot\operatorname{pr}_{1,2}^{*}X_{\phi}$ $\operatorname{pr}_{3,*}(z\cdot\operatorname{pr}_{1,2}^{*}X_{\phi})$
$\Delta^{3}_{C}$ $\Delta^{3}_{C}(F_{\phi})$ $F_{\phi}$
$-\Delta_{C}\times e$ $-\Delta^{2}_{C}(F_{\phi})\times e$ $-\deg(F_{\phi})e$
$-\operatorname{pr}_{1,3}^{*}\Delta_{C}\cdot\operatorname{pr}_{2}^{*}e$ $-\operatorname{pr}_{1,3}^{*}\Delta_{C}(\phi^{\vee}(e))\cdot\operatorname{pr}_{2}^{*}e$ $-\phi^{\vee}(e)$
$-e\times\Delta_{C}$ $-e\times\Delta_{C}(\phi(e))$ $-\phi(e)$
$C\times e\times e$ $\phi^{\vee}(e)\times e$ $\deg(\phi^{\vee})e$
$e\times C\times e$ $e\times\phi(e)\times e$ $\deg(\phi)e$
$e\times e\times C$ 00 00
Table 1. Components of the calculation of $\Gamma^{3}(C,e)_{*}X_{\phi}$
Combining these terms, we get the following point, $F_{\phi}-\deg(F_{\phi})e-\phi(e)-\phi^{\vee}(e)+\deg(\phi)e+\deg(\phi^{\vee})e$, which is commonly the definition of a Chow-Heegner point. Multiplying by $22g-$ 2 to clear denominators gives $\operatorname{Sh}(\phi)=(2g-2)F-\deg(F)K_{C}-\phi(K_{C})-\phi^{\vee}(K_{C})+(\deg(\phi)+\deg(\phi^{\vee}))K_{C}$. If the Ceresa cycle is zero in the rational Chow group, then so is $\Gamma^{3}(C,e)$, and so pushing forward $X_{\phi}$ through $\Gamma^{3}(C,e)$ must also be zero. This shows the resulting divisor is trivial in the Jacobian tensored with $\mathbb{Q}$, and so it is torsion. ∎
2.3. Fixed Points of Hecke Operators
In the case of modular curves there are many non-trivial endomorphisms of the Jacobian, namely Hecke operators. More precisely, for $C=X_{0}(N)$ the endomorphism ring of the Jacobian contains the Hecke algebra generated (over $\mathbb{Z}$) by the induced Atkin-Lehener operators $w_{d}$ for all $d|N$ and coprime to $\frac{N}{d}$; and for all $\ell\nmid N$ the Hecke operator $T_{\ell}$.
As we now explain, the set of fixed points of Hecke operators contain CM points on the $X_{0}(N)$, that is, non-cuspidal points parametrising an elliptic curve with complex multiplication. For ease of exposition, we introduce the following notation.
Notation.
For any integer $d\geq 1$, we write $D_{d}$ for the effective divisor on $X_{0}(N)$, supported on CM points corresponding to an order with discriminant $-d$ and multiplicity $1$. If no such points exist on $X_{0}(N)$, we take $D_{d}$ to be the zero divisor.
For $\phi=w_{d}$, $\phi=\phi^{\vee}$ and the degrees of the correspondences are 1. This gives the following expression for the shadow, since the canonical is invariant under automorphisms.
(1) $\operatorname{Sh}(\phi)=(2g-2)F_{\phi}-\deg(F)K_{C}.$
The fixed points of Atkin-Lehner operators are well-know due to Ogg [31, Proposition 3].
Proposition 5.
The points in the support of $F_{w_{d}}$ are either:
  • cusps if $d=4$;
  • CM points with endomorphism ring containing $\mathbb{Z}[\sqrt{-d}]$
Moreover, all fixed points occur with multiplicity $1$.
In particular, for prime level $N=p$, the fixed point divisor for $\phi=w_{p}$ is simply $D_{p}+D_{4p}$, where we recall our convention that if $p$ is not the discriminant of a CM order $D_{p}=0$.
The second type of Hecke operator $T_{\ell}$ where $\ell\nmid N$, can be described via a square as follows, where the vertical arrows are both the forgetful map
${X_{0}(N\ell)}$${X_{0}(N\ell)}$${X_{0}(N)}$${X_{0}(N)}$$\scriptstyle{w_{\ell}}$
see [13, Chapter 5] for details. As a result of this, we see that in the sense of the previous section, $T_{\ell}=T_{\ell}^{\vee}$. Further, as is well known, the degrees for both are $\ell+1$. This simplifies $\operatorname{Sh}(T_{\ell})$ to $(2g-2)F_{T_{\ell}}-\deg(F_{T_{\ell}})K_{X_{0}(N)}-2T_{\ell}(K_{X_{0}(N)})+2(\ell+1)K_{X_{0}(N)}$.
Proposition 6.
The points in the support of $F_{T_{\ell}}$ are either:
  • Cusps;
  • CM elliptic curves, with endomorphism ring $\mathcal{O}$ containing an element of norm $\ell$, and an ideal, $I$ such that $\mathcal{O}/I$ is cyclic of order $N$.
Proof.
The cusps are fixed as there are 2 cusps of $X_{0}(N\ell)$ lying above each cusp of $X_{0}(N)$, and the Atkin-Lehner operator $w_{\ell}$ interchanges them.
Now suppose $(E,C_{N})$ is a fixed point of the Hecke operator, then there exists an order $\ell$ subgroup, $C_{\ell}$, such that $(E,C_{N})\cong(E/C_{\ell},(C_{N}+C_{\ell})/C_{\ell})$. In particular, $E$ has an endomorphism of order $\ell$. The last condition is imposing the existence of a suitable order $N$ subgroup. ∎
It remains to compute the multiplicities with which these occur. There are two reasons for this multiplicity to be greater than one, the first is that the graph of $X_{0}(N\ell)$ in $X_{0}(N)^{2}$ has a node at this point, the second is that the graph is tangent to the diagonal here. We first exclude the second possibility.
Proposition 7.
Suppose $P\in X_{0}(N\ell)$ has both projections to $X_{0}(N)$ equal, then the image of the tangent space of $P$ under the map from $X_{0}(N\ell)$ to $X_{0}(N)^{2}$ is not diagonal, for $\ell=2,3,5$.
Proof.
We start with the case that $P$ is a cusp. In this case, the local parameter at $P$ is either the same as for its image in $X_{0}(N)$, or its $\ell$-th root, and the opposite is true for $w_{\ell}(P)$. In either case, the projection of its tangent space to $X_{0}(N)^{2}$ is along one of the natural fibrations.
Otherwise, $P$ corresponds to an elliptic curve. By the previous proposition, this elliptic curve has CM. Since the $j$-invariant is a local parameter for CM points on $X_{0}(N)$, it is enough to check that $j-j\circ\omega_{\ell}$ has simple zeros at CM points on $X_{0}(N\ell)$. As this function is a pullback from $X_{0}(\ell)$, we start by proving this for $X_{0}(\ell)$. For $\ell=2$, $j-j\circ\omega_{\ell}$ has two poles of order 2, and so is a degree 4 function. There are 3 CM orders with an element of norm 2, $\mathbb{Z}[\sqrt{-1}]$, $\mathbb{Z}[\sqrt{-2}]$ and $\mathbb{Z}[\frac{-1+\sqrt{-7}}{2}]$. The first two of these have a unique ideal of index 2, and the second has 2 choices, giving rise to a total of 4 points on $X_{0}(2)$ where this function vanishes. This forces all 4 points to be simple zeros. Similar counting arguments hold for $\ell=3,5$, where for $\ell=5$, class numbers also have to be taken into account. It remains to show that the simple zeros on $X_{0}(\ell)$ pullback to simple zeros on $X_{0}(N\ell)$. This can only fail if $X_{0}(N\ell)\to X_{0}(\ell)$ ramifies over these points, but this does not happen (e.g. by considering the order of vanishing of $j$). ∎
This result should hold more generally by considering fist order deformations of lattices, however, as we only use these 3 cases, we simply prove it for these, mainly for exposition reasons.
We now handle the other source of higher multiplicity, where multiple distinct points of $X_{0}(N\ell)$ map to the same point in $X_{0}(N)^{2}$.
Proposition 8.
The multiplicity of cusps in $F_{T_{\ell}}$ are 2, and for an elliptic curve with CM by $\mathcal{O}$, it is the number of prime ideals above $\ell$ in $\mathcal{O}$.
Proof.
The multiplicity is determined by number of points of $X_{0}(N\ell)$ above each point of the diagonal in $X_{0}(N)^{2}$.
For each cusp of $X_{0}(N)$, there are exactly 2 cusps of $X_{0}(N\ell)$ above it, and both map to the same point in $X_{0}(N)^{2}$, namely the diagonal embedding of the cusp.
For a point of $X_{0}(N\ell)$ corresponding to a CM elliptic curve, the points above it of interest, are the ones where the $\ell$-structure comes from the CM structure, and so the $\ell$-structure comes from a choice of ideal of index $\ell$. Any such choice is appropriate, as if one ideal is principal, so is the other. ∎
Combining the above in the special case of $T_{2}$, gives the following
$F_{T_{2}}=D_{4}+D_{8}+2D_{7}+2D_{\infty}$
where $D_{\infty}$ is the cuspidal divisor (all with multiplicity $1$) and we highlight the fact that if no such CM point exists, $D_{d}$ is assumed to be zero.
3. The Arithmetic and Geometry of Modular Curves
3.1. Torsion Subgroups of Modular Jacobians
Throughout this section $p\geq 5$ denotes a prime number. The proof of Theorem 1 focuses on proving that the shadow $D$ of some Hecke operator on $J_{0}(p)$ is non-torsion. The rational torsion subgroup of $J_{0}(p)$ is completely understood due to the classical work of Mazur and Ogg. In this section we give an overview of some well known results. Note that $X_{0}(p)$ has two cusps, both rational, which we denote by $c_{0}$ and $c_{\infty}$.
We first recall the following proposition due to Ogg [30].
Proposition 9 (Ogg).
The linear equivalence class of $c_{0}-c_{\infty}$ in $J_{0}(p)$ has order $\frac{p-1}{\text{gcd}\left(p-1,12\right)}$.
In the same work, Ogg conjectured that the linear equivalence of this divisor generates the entire rational torsion subgroup of $J_{0}(p)$. This was proved by Mazur using the theory of Eisenstein ideals [29, Page 142].
Theorem 10 (Mazur).
$J_{0}(p)\left(\mathbb{Q}\right)_{\text{tors}}=\langle[c_{0}-c_{\infty}]\rangle.$
To prove that $D\in J_{0}(p)\left(\mathbb{Q}\right)$ is non-torsion, we further project to the quotient of $J_{0}\left(p\right)$ by the Atkin-Lehner involution $w_{p}$, that is the involution on $J_{0}(p)$ induced by the Atkin-Lehner involution on the curve. We write $X_{0}^{+}(p)$ for the quotient curve and $J_{0}^{+}(p)$ for the quotient of the Jacobian. In the same reference, Mazur proved the following.
Theorem 11 (Mazur).
If $X_{0}^{+}(p)$ has positive genus, that is $p>73$ or $p=37,43,53,61,67$, then $J_{0}^{+}(p)\left(\mathbb{Q}\right)_{\text{tors}}$ is trivial.
The proof makes us of the following results, which also appear in some of our arguments.
Proposition 12 (Mazur, Lemma III.2.6).
Let $d$ be an integer. If $H^{0}\left(X_{0}^{+}\left(p\right),\mathcal{O}\left(dc_{\infty}\right)\right)$ is non-trivial, then $d>\frac{p}{96}$.
Proposition 13 (Mazur, Proposition III.2.8).
Let $\mathcal{O}$ be an order in a quadratic imaginary field. Then
$h\left(\mathcal{O}\right)\leq\frac{1}{3}|\Delta|^{1/2}\log|\Delta|$
where $\Delta$ is the discriminant of $\mathcal{O}$.
3.2. Canonical Divisors
As the canonical embedding is used throughout our study of Ceresa cycles, we determine expressions for the canonical divisor of modular curves.
Proposition 14.
The canonical divisor of the modular curve $X_{0}(p)$ is given as follows.
  • If $p\equiv 11\mod{12}$, $K_{X_{0}(p)}=\frac{p-11}{12}(c_{0}+c_{\infty})$
  • If $p\equiv 7\mod{12}$, $3K_{X_{0}(p)}=\frac{p-11}{4}(c_{0}+c_{\infty})-2D_{3}$
  • If $p\equiv 5\mod{12}$, $2K_{X_{0}(p)}=\frac{p-11}{6}(c_{0}+c_{\infty})-D_{4}$
  • If $p\equiv 1\mod{12}$, $6K_{X_{0}(p)}=\frac{p-11}{2}(c_{0}+c_{\infty})-4D_{3}-3D_{4}$
Proof.
Consider the $j$-map, $j:X_{0}(p)\to\mathbb{P}^{1}$. By the Riemann-Hurwitz formula, $K_{X_{0}(p)}=j^{*}(K_{\mathbb{P}^{1}})+R$, where $R$ is the ramification divisor. The canonical divisor of $\mathbb{P}^{1}$ is twice a point, and we can take that point to be $\infty$. The ramification divisor is $(p-1)c_{\infty}$, along with the points with $j=0$ (except where the non-trivial automorphism stabilises the $p$-subgroup) with multiplicity 2, and the points with $j=1728$ with the same proviso and multiplicity 1. This shows the canonical divisor on $X_{0}(p)$ is $-2pc_{\infty}-2c_{0}+(p-1)c_{\infty}+2\{j=0\}-2D_{3}+\{j=1728\}-D_{4}$. Since $3\{j=0\}-2D_{3}$ is the vanishing of $j$, this divisor is linearly equivalent to $pc_{\infty}+c_{0}$, and similarly $2\{j=1728\}-D_{4}\sim pc_{\infty}+c_{0}$. We deduce $6K_{X_{0}(p)}=-6(p+1)c_{\infty}-12c_{0}+4pc_{\infty}+4c_{0}-4D_{3}+3pc_{\infty}+3c_{0}-3D_{4}=(p-6)c_{\infty}-5c_{0}-4D_{3}-3D_{4}$. Using $\frac{p-1}{2}c_{\infty}\sim\frac{p-1}{2}c_{0}$, this takes the more symmetric form $\frac{p-11}{2}(c_{\infty}+c_{0})-8D_{3}-3D_{4}$.
If $p\equiv 11\mod{12}$, then $D_{3}=D_{4}=0$. In this case $6K_{X_{0}(p)}=\frac{p-11}{2}(c_{\infty}+c_{0})$. We now consider $K_{X_{0}(p)}-\frac{p-11}{12}(c_{\infty}+c_{0})$. This is a rational divisor, and has order dividing 6 in the Jacobian. By Ogg’s conjecture, proven by Mazur, the rational torsion has order dividing the numerator of $\frac{p-1}{12}$. By the congruence condition on $p$, the numerator is coprime to 6. This forces $K_{X_{0}(p)}-\frac{p-11}{12}(c_{\infty}+c_{0})$ to have order 1, and so equality holds.
The other cases follow similarly. ∎
A similar result holds for the canonical divisors of $X_{0}(N)$, with an identical proof.
Proposition 15.
For any integer $N$, the canonical divisor satisfies $6K_{X_{0}(N)}=D^{\prime}-4D_{3}-3D_{4}$, for some divisor $D^{\prime}$ supported on the cusps, and $D_{3}$ and $D_{4}$ are understood to be zero when there are no points with CM by those discriminants
3.3. The Minimal Regular Models of $X_{0}(p)$ and $X_{0}^{+}(p)$
A key ingredient of the proof of Theorem 1 is reduction modulo $p$. We give a brief description of the special fibres of the minimal regular models of $X_{0}(p)$ and $X_{0}^{+}(p)$ over $\mathbb{F}_{p}$.
The reductions of $X_{0}\left(p\right)$ and $X_{0}^{+}\left(p\right)$ are smooth away from $p$. The special fibre of the minimal regular model of $X_{0}(N)$ at $p|N$, where $N$ is square-free and coprime to $6$, is described in [12] and [29, Appendix]. We recall this for prime level $N=p\geq 5$
Theorem 16.
The special fibre of the minimal regular model of $X_{0}\left(p\right)_{/\mathbb{Z}_{p}}$ consists of two copies of /$\mathbb{P}^{1}_{/\mathbb{\overline{F}}_{p}}$, each with multiplicity $1$, crossing transversely at the supersingular points. More precisely, the crossing points are non-cuspidal points parametrising supersingular elliptic curves, with a specified subgroup of order $p$. There are precisely $k$ crossing points corresponding to supersingular elliptic curves with $j\neq 0,1728$, where
$k=\frac{p-a}{12}$ where $p\equiv a\mod{12}$.
If $j\in\{0,1728\}$ is a supersingular $j$-invariant, we replace the crossing by one projective line if $j=1728$ and by a chain of two lines if $j=0$, connecting the two components.
Recall that $j=0$ is supersingular if and only if $p\equiv 1\mod{6}$, and $j=1728$ is supersingular if and only if $p\equiv 1\mod{4}$.
The minimal regular model of $X_{0}^{+}\left(p\right)$ is obtained from that of $X_{0}\left(p\right)$ as described by Xue [37].
Theorem 17.
The special fibre of the minimal regular model of $X_{0}^{+}\left(p\right)_{/\mathbb{Z}_{p}}$ over $\mathbb{F}_{p}$ is a rational curve with $\frac{1}{2}S_{p^{2}}$ nodes where $S_{p^{2}}$ is the number of supersingular $j$-invariants that are not in $\mathbb{F}_{p}$.
4. Proof of Theorem 1
Throughout this section we write $p\geq 5$ for a prime number, $X_{0}(p)$ for the complete modular curve associated to the congruence subgroup $\Gamma_{0}(p)$ and $X=X_{0}^{+}(p)$ for the quotient curve obtained from the unique Atkin-Lehner involution $w_{p}$. We also write $w_{p}$ for the induced involution on the Jacobian $J_{0}(p)$ of $X_{0}(p)$ and $J=J_{0}^{+}(p)$ for the quotient variety. In this case we are able to identify $J$ with the Jacobian of $X$ since the map quotient map $\pi:X_{0}(p)\rightarrow X$ is ramified.
The proof of Theorem 1 reduces to proving that the scaled shadow of some carefully chosen Hecke operator defines a point of infinite order in the Jacobian. We consider various points, principally $\operatorname{Sh}(T_{2})$, but also $\operatorname{Sh}(T_{3})$ and $\operatorname{Sh}(w_{p})$ depending on the congruence class of $p$. We denote this point by $D$ and note the following trivial corollary of Theorems 10 and 11.
Corollary 18.
Suppose $[\pi(D)]\in J(\mathbb{Q})$ is non-zero. Then $[D]\in J_{0}(p)(\mathbb{Q})$ is of infinite order.
4.1. Reduction Arguments
For most congruence classes we map the shadow point to $J$ and reduce modulo $p$ to conclude that it is non-zero. In particular, we use the following results.
Proposition 19.
There exists a homomorphism
$\mathrm{red}:J(\mathbb{Q}_{p})\longrightarrow\operatorname{Pic}^{0}(\mathcal{X}_{/\mathbb{F}_{p}})$
where $\mathcal{X}_{/\mathbb{F}_{p}}$ is the special fibre of the minimal regular model of $X_{/\mathbb{\mathbb{Q}}_{p}}$.
Proof.
Let $\mathcal{J}/\mathbb{Z}_{p}$ be the Néron model of $J_{/\mathbb{Q}_{p}}$ and write $\mathcal{J}_{/\mathbb{F}_{p}}$ for its special fibre. By the Néron mapping property
$J(\mathbb{Q}_{p})\cong\mathcal{J}(\mathbb{Z}_{p})$.
We write $\mathcal{J}^{0}$ for the connected component of the identity and $\Phi$ for the component group. By definition
$\mathcal{J}/\mathcal{J}^{0}\cong\Phi$,
and since $J$ is the Jacobian of a curve, we can determine the component group from the special fibre of the minimal regular model of the curve, see [5, Section 9.6, Theorem 1]. As the special fibre described in Theorem 17 has a single component we deduce that $\Phi$ is trivial, and hence $\mathcal{J}(\mathbb{F}_{p})=\mathcal{J}^{0}(\mathbb{F}_{p})$. Moreover, by [5, Section 9.5, Theorem 1], $\mathcal{J}^{0}(\mathbb{F}_{p})\cong\operatorname{Pic}^{0}(\mathcal{X}_{/\mathbb{F}_{p}})$ and $\mathrm{red}$ is simply the composition
$J(\mathbb{Q}_{p})\xrightarrow{\sim}\mathcal{J}(\mathbb{Z}_{p})\rightarrow\mathcal{J}(\mathbb{F}_{p})=\mathcal{J}^{0}(\mathbb{F}_{p})\xrightarrow{\sim}\operatorname{Pic}^{0}(\mathcal{X}_{/\mathbb{F}_{p}})$.
We now give a criterion for a divisor to be non-vanishing in $\operatorname{Pic}^{0}(\mathcal{X}/\mathbb{F}_{p})$.
Proposition 20.
Let $D=\sum_{i}n_{i}a_{i}$ be a divisor on $\mathcal{X}/\mathbb{F}_{p}$, and define $f(x)=\prod_{i}(x-a_{i})^{n_{i}}$, by identifying the special fibre of $X$ with a $\mathbb{P}^{1}$. If $j\in\mathbb{F}_{p^{2}}\setminus\mathbb{F}_{p}$ is a supersingular $j$-invariant such that $f(j)\notin\mathbb{F}_{p}$, then $D\neq 0$.
Proof.
As $j$ is a supersingular $j$-invariant, in $\mathcal{X}/\mathbb{F}_{p}$ it is glued to $\bar{j}$, where $\overline{}$ denotes the action of Frobenius. Since $f(j)\notin\mathbb{F}_{p}$, then $f(j)\neq\overline{f(j)}=f\left(\overline{j}\right)$ In particular, $f$ is not a well-defined function on $\mathcal{X}/\mathbb{F}_{p}$, and so $D$ is not trivialised by it. ∎
The next proposition describes one way of applying the previous proposition in characteristic 0.
Proposition 21.
Let $D=\sum_{i}n_{i}P_{i}$ be a rational, degree 0 divisor on $X_{0}(p)$. Take a CM order, $\mathcal{O}$ such that $p$ is inert in both $\mathcal{O}$ and the Hilbert class field of $\mathcal{O}$. Define $\alpha=\prod_{i}(j_{\mathcal{O}}-j(P_{i}))^{n_{i}}$, where $j_{\mathcal{O}}$ is a $j$-invariant attached to $\mathcal{O}$. If there does not exist an $a\in\mathbb{Z}$ such that the $\mathfrak{p}$-adic valuation of $\alpha-a$ is strictly positive for some $\mathfrak{p}$ lying over $p$, $\pi(D)$ is non-zero in $J$.
Proof.
We first push $D$ forwards under $\pi$, to obtain an element of $J$. By Proposition 19, we have a reduction map $J\to\operatorname{Pic}^{0}(\mathcal{X}/\mathbb{F}_{p})$. Under this map, $D$ is sent to $\sum_{i}n_{i}j(P_{i})$. Due to the assumptions on $\mathcal{O}$, $j_{\mathcal{O}}$ is a supersingular $j$-invariant mod $\mathfrak{p}$, which is not in $\mathbb{F}_{p}$. As $\alpha=f(j_{\mathcal{O}})$, where $f$ is as in the previous proposition, the assumption on the $\mathfrak{p}$-adic valuation translates into $f(j_{\mathcal{O}})\notin\mathbb{F}_{p}$, and so the reduction of $\pi(D)$ is non-zero. ∎
4.2. Primes congruent to 5 or 7 modulo 12
We work with the divisor $D=3\cdot\operatorname{Sh}(T_{2})$ or $6\cdot\operatorname{Sh}(T_{2})$ depending on the canonical divisor and apply Proposition 21 to $D$. The explicit dependence of the coefficients on $D$ pose an obstacle, as otherwise the quantity $\alpha$ could be computed without use of $p$, for some choice of $\mathcal{O}$ which works for the whole congruence class. We circumvent this issue by observing that as $X$ has non-split semistable reduction, the group $\operatorname{Pic}^{0}(\mathcal{X}/\mathbb{F}_{p})$ has exponent $p+1$, and so we may reduce the coefficients of $D$ modulo $p+1$ for the purposes of this calculation.
For each congruence class of $p$ within these classes, we can now compute the value $\alpha$, and determine which primes do not have the necessary valuation condition. This can be achieved by factorising $\alpha-\sigma(\alpha)$, where $\sigma$ is a Frobenius element of the Galois group for $p$. In practice, this factorisation can be unwieldy. We avoid this by using 2 CM orders, and taking the greatest common divisor of the two values, and then factorising. In Table 2, we list the divisors $D$, and in Table 3 we list the CM orders with the corresponding $j$-invariants used, along with any primes where Proposition 21 does not apply. The exceptions can all be seen to be primes where $X_{0}(p)$ is hyperelliptic. These were computed using the jTest routine in divisor_checker.m.
$p$ $D$
$p\equiv 7\mod{12}$, $p\equiv 7\mod{8}$, $p\equiv 3,5,6\mod{7}$ $3\operatorname{Sh}(T_{2})=12D_{12}-4D_{3}-8(c_{0}+c_{\infty})$
$p\equiv 7\mod{12}$, $p\equiv 7\mod{8}$, $p\equiv 1,2,4\mod{7}$ $3\operatorname{Sh}(T_{2})=\left(p-19\right)D_{7}+12D_{12}+4D_{3}+(-p+3)\left(c_{0}+c_{\infty}\right)$
$p\equiv 7\mod{12}$, $p\equiv 3\mod{8}$, $p\equiv 3,5,6\mod{7}$ $6\operatorname{Sh}(T_{2})=\left(p-19\right)D_{8}+24D_{12}+(-p-5)(c_{0}+c_{\infty})$
$p\equiv 7\mod{12}$, $p\equiv 3\mod{8}$, $p\equiv 1,2,4\mod{7}$ $6\operatorname{Sh}(T_{2})=(p-19)D_{8}+(2p-38)D_{7}+16D_{3}+24D_{12}+(-3p+17)(c_{0}+c_{\infty})$
$p\equiv 5\mod{12}$, $p\equiv 5\mod{8}$, $p\equiv 3,5,6\mod{7}$ $6\operatorname{Sh}(T_{2})=(p-11)D_{4}+12D_{16}+(-p-1)(c_{0}+c_{\infty})$
$p\equiv 5\mod{12}$, $p\equiv 5\mod{8}$, $p\equiv 1,2,4\mod{7}$ $6\operatorname{Sh}(T_{2})=(p+1)D_{4}+2(p-17)D_{7}+12D_{16}+(-3p+21)(c_{0}+c_{\infty})$
$p\equiv 5\mod{12}$, $p\equiv 1\mod{8}$, $p\equiv 3,5,6\mod{7}$ $6\operatorname{Sh}(T_{2})=(p-5)D_{4}+(p-17)D_{8}+12D_{16}+(-2p+10)(c_{0}+c_{\infty})$
$p\equiv 5\mod{12}$, $p\equiv 1\mod{8}$, $p\equiv 1,2,4\mod{7}$ $6\delta_{T_{2}}=(p+7)D_{4}+2(p-17)D_{7}+(p-17)D_{8}+12D_{16}+(-4p+32)(c_{0}+c_{\infty})$
$p\equiv 1\mod{12}$, $p\equiv 5\mod{8}$, $p\equiv 3,5,6\mod{7}$ $6\operatorname{Sh}(T_{2})=(p-19)D_{4}+12D_{16}+24D_{12}+(-p-17)(c_{0}+c_{\infty})$
$p\equiv 1\mod{12}$, $p\equiv 5\mod{8}$, $p\equiv 1,2,4\mod{7}$ $6\operatorname{Sh}(T_{2})=(p-7)D_{4}+2(p-25)D_{7}+16D_{3}+12D_{16}+24D_{12}+(-3p+5)(c_{0}+c_{\infty})$
$p\equiv 1\mod{12}$, $p\equiv 1\mod{8}$, $p\equiv 3,5,6\mod{7}$ $6\operatorname{Sh}(T_{2})=(p-13)D_{4}+(p-25)D_{8}+8D_{3}+24D_{12}+12D_{16}+(-2p-6)(c_{0}+c_{\infty})$
$p\equiv 1\mod{12}$, $p\equiv 1\mod{8}$, $p\equiv 1,2,4\mod{7}$ $6\operatorname{Sh}(T_{2})=24D_{3}+(p-1)D_{4}+2(p-25)D_{7}+(p-25)D_{8}+24D_{12}+12D_{16}+(-4p+16)(c_{0}+c_{\infty})$
Table 2. Scaled Shadow Points
$p$ $j$ $\Delta_{j}$ Exceptions
$p\equiv 7\mod{12}$, $p\equiv 7\mod{8}$, $p\equiv 3,5,6\mod{7}$ $76771008+44330496\sqrt{3}$ $-36$ $7,31$
$p\equiv 7\mod{12}$, $p\equiv 7\mod{8}$, $p\equiv 1,2,4\mod{7}$ $76771008+44330496\sqrt{3}$ $-36$emptyset
$p\equiv 7\mod{12}$, $p\equiv 3\mod{8}$, $p\equiv 3,5,6\mod{7}$ $76771008+44330496\sqrt{3}$ $-36$ $19$
$p\equiv 7\mod{12}$, $p\equiv 3\mod{8}$, $p\equiv 1,2,4\mod{7}$ $76771008+44330496\sqrt{3}$, $41113158120+29071392966\sqrt{2}$ $-36$, $-64$emptyset
$p\equiv 5\mod{12}$, $p\equiv 5\mod{8}$, $p\equiv 3,5,6\mod{7}$ $1417905000+818626500\sqrt{3}$ $-48$ 5
$p\equiv 5\mod{12}$, $p\equiv 5\mod{8}$, $p\equiv 1,2,4\mod{7}$ $1417905000+818626500\sqrt{3}$, $26125000+18473000\sqrt{2}$ $-48$, $-32$ $29$
$p\equiv 5\mod{12}$, $p\equiv 1\mod{8}$, $p\equiv 3,5,6\mod{7}$ $1417905000+818626500\sqrt{3}$, $137458661985000-51954490735875\sqrt{7}$ $-48$, $-112$ $17$, $41$
[…]
$p\equiv 1\mod{12}$, $p\equiv 5\mod{8}$, $p\equiv 3,5,6\mod{7}$ $26125000+18473000\sqrt{2}$ $-32$ $13$
$p\equiv 1\mod{12}$, $p\equiv 5\mod{8}$, $p\equiv 1,2,4\mod{7}$ $26125000+18473000\sqrt{2}$, $188837384000+77092288000\sqrt{6}$ $-32$, $-72$ $37$
[…]
Table 3. Supersingular $j-$invariants.
4.3. Primes congruent to 1 modulo 12
The methods of the previous section continue to apply for $p\equiv 1\mod{12}$, except for the class $p\equiv 1\mod{8}$ and $p\equiv 1,2,4\mod{7}$. For primes in this class, there is no longer a uniform choice of CM order with the necessary properties. The shadows are listed in Table 2, and for the other 3 congruence classes, the choice of $\mathcal{O}$ and $j$ are in Table 3.
To handle the remaining cases, we show that at least one of $\operatorname{Sh}(T_{2})$ and $\operatorname{Sh}(T_{3})$ are of infinite order.
Let $f_{1}$ and $f_{2}$ be the functions attached to the pushforwards of $6\operatorname{Sh}(T_{2})$ and $6\operatorname{Sh}(T_{3})$ respectively, as in Proposition 20. As before, we reduce the coefficients of $6\operatorname{Sh}(T_{2})$ and of $6\operatorname{Sh}(T_{3})$ modulo $p+1$, as this does not affect their reduction. Define the curve $\tilde{C}_{i}$ to be fibre product $\mathbb{P}^{1}\times_{f_{i}}\mathbb{P}^{1}$ less the diagonal. Further, let $C_{i}$ be the quotient of $\tilde{C}_{i}$ under the involution swapping the factors and assume $C_{i}$ is embedded in $\mathbb{P}^{2}=\mathrm{Sym}^{2}\mathbb{P}^{1}$.
Proposition 22.
Assume that the pushforwards of $6\operatorname{Sh}(T_{2})$ and $6\operatorname{Sh}(T_{3})$ are both zero under the reduction map in Proposition 19. Then #$\#C_{1}(\mathbb{F}_{p})\cap C_{2}(\mathbb{F}_{p})\geq g(X_{0}^{+}(p))$, with notation as above.
Proof.
As the pushforward of $6\operatorname{Sh}(T_{2})$ is zero under the reduction map, for every supersingular $j$-invariant, $f_{1}(j)\in\mathbb{F}_{p}$ by Proposition 20. In particular, $f_{1}(j)=f_{1}(\overline{j})$, and so $(j,\overline{j})$ is a point on the fibre product. If $j$ is not defined over $\mathbb{F}_{p}$, this point is not on the diagonal, and so defines a point on $\tilde{C}_{1}$. This reduces to a point on $C_{1}$, and as $j$ is defined over $\mathbb{F}_{p^{2}}$, it becomes rational. The same argument applies for $f_{2}$ and so this point appears on both curves. The bound in the statement of the proposition then follows as there are $g(X_{0}^{+}(p))$ such pairs of $j$-invariants. ∎
Unless $C_{1}$ and $C_{2}$ share a common component, the size of the intersection is bounded by Bezout’s theorem. We therefore study the components of $C_{1}$.
Proposition 23.
All components of $C_{1}$ contain the image of the point $(1728,8000)\in\tilde{C}_{1}$.
Proof.
From Table 2, we see that
$f_{1}=\frac{x^{24}(x-54000)^{12}(x-287496)^{12}}{(x-1728)^{2}(x-8000)^{26}(x+3375)^{52}}.$
This is a perfect square, so let $f_{1}=g_{1}^{2}$. This forces $\tilde{C}_{1}$ to have at least 2 components, one where $g_{1}$ agrees on the two points, and the other where $g_{1}$ takes opposite signs. These are stable under swapping the order of points and so $C_{1}$ has at least 2 components.
Over $\mathbb{Q}$, $\tilde{C}_{1}$ has exactly 2 irreducible (but not necessarily geometrically irreducible) components by direct calculation in Magma, which must be the two already described. The point $(1728,8000)$ is on both, since it is a pole for $g_{1}$. As the point is rational, it will be on all geometrically irreducible components comprising the irreducible components.
It now remains to determine when these geometrically irreducible components split further modulo $p$. The number of such components is related to the Galois group of $f_{1}$. Namely, let $G$ be the geometric Galois group, and it acts naturally on a set of order 80. The number of geometrically irreducible components of $C_{1}$ is the number of orbits of $G$ acting on unordered pairs from the set of 80 elements, and so it is enough to understand how the geometric Galois group behaves under reduction. By Lemma 2 in [3], this can be deduced from the monodromy information. In particular, if $p$ does not divide the order of any monodromy element, and no ramification points above a fixed branch point collide modulo $p$, then reduction is injective on the Galois group. There are 7 branch points, 00, $\infty$ and a degree 5 orbit. We first check 00, and $f_{1}^{-1}(0)$ consists of 4 points, 00, $54000$, $287496$ and $\infty$. The differences of these are not divisible by any prime congruent to 1 mod 12. Similarly, no difference between the points over $\infty$ are divisible by a prime congruent to 1 mod 24. Finally, if there were a collision above any of the other 5 branch points, this would violate the Riemann-Hurwitz formula for the genus, unless two of the branch points also collided. Analysing the polynomial shows that again, this cannot happen for a suitable $p$.
We therefore conclude that all irreducible components of $C_{1}$ over $\overline{\mathbb{Q}}$ remain irreducible modulo $p$, and so all irreducible components modulo $p$ contain the given point. ∎
The value of $6\operatorname{Sh}(T_{3})$ depends on the congruence class of $p\mod{11}$. If $p$ is a non-residue modulo 11,
$6\operatorname{Sh}(T_{3})=(p-1)D_{3}+12D_{4}+2(p-25)D_{8}+(p-25)D_{12}+24D_{27}+12D_{36}-4(p-4)(c_{\infty}+c_{0}),$
and
$f_{2}=\frac{(x-1728)^{12}(x+12288000)^{24}(x^{2}-153542016x-1790957481984)^{12}}{x^{2}(x-8000)^{52}(x-54000)^{26}}$
If $p$ is a residue,
$6\operatorname{Sh}(T_{3})=(p+15)D_{3}+24D_{4}+2(p-25)D_{8}+2(p-25)D_{11}+(p-25)D_{12}+24D_{27}+12D_{36}-(6p-38)(c_{\infty}+c_{0})$
and
$f_{2}=\frac{x^{12}(x-1728)^{24}(x+12288000)^{24}(x^{2}-153542016x-1790957481984)^{12}}{(x-8000)^{52}(x+32768)^{52}(x-54000)^{26}}.$
In either case, we see that the image of $(1728,8000)$ cannot be on $C_{2}$, since 1728 is a zero of $f_{2}$ in both cases, and 8000 is a pole. In particular, $C_{2}$ cannot contain a component of $C_{1}$, since all components of $C_{1}$ contain the image of $(1728,8000)$. We can now apply Bezout’s theorem. If $p$ is not a residue modulo 11, then the degree of $C_{1}$ and $C_{2}$ are both 79 (the degree of $f_{1}$ and $f_{2}$, minus 1 for the diagonal component that has been removed), and so #$\#C_{1}\cap C_{2}\leq 79^{2}$. Similarly, if $p$ is a residue, then #$\#C_{1}\cap C_{2}\leq 79\times 129$. Combining this inequalities gives an upper bound for the genus of any curve with both $\operatorname{Sh}(T_{2})$ and $\operatorname{Sh}(T_{3})$ torsion.
As $g(X_{0}^{+}(p))=\frac{p-1}{24}-\frac{1}{4}h(-4p)$, any $p$ with both divisors torsion has $h(-4p)\geq\frac{p-1}{6}-4\times 79^{2}$ if $p$ is a non-residue modulo 11, and $h(-4p)\geq\frac{p-1}{6}-4\times 79\times 129$ otherwise. Applying the class number bound, Proposition 13, gives $2\sqrt{p}\log(4p)\geq\frac{p-1}{2}-12\times 79^{2}$, which gives $p\leq 172090$. There are 430 primes in this congruence class where genus is below the bound, and the largest is 150889. For each prime, we check $6\operatorname{Sh}(T_{2})\neq 0$ using Proposition 20.
Similarly, if $p$ is a residue, the bound is $p\leq 273684$, and there are 642 primes in this class with genus below the bound, and of these, 244897 is the largest. Again, applying Proposition 20 shows that for all of them, $6\operatorname{Sh}(T_{2})$ is non-zero. These cases are checked in the file primes1T2.m.
4.4. Primes congruent to 11 modulo 12
When $p\equiv 11\mod{12}$ we cannot guarantee that $\operatorname{Sh}(T_{2})\neq 0$, or indeed $\operatorname{Sh}(T_{\ell})$ for any fixed prime $\ell$. We instead consider $\operatorname{Sh}(w_{p})$,
$\operatorname{Sh}(w_{p})=\left(g-1\right)\left(D_{p}+D_{4p}\right)-n(g-1)\left(c_{0}+c_{\infty}\right)$
where $n=\frac{1}{2}\left(h\left(-p\right)+h\left(-4p\right)\right)$, and $g$ is the genus of $X_{0}(p)$.
If $[\operatorname{Sh}(w_{p})]]$ is torsion in $J_{0}(p)(\mathbb{Q})$, we divide by $g-1$ and project onto $X_{0}^{+}(p)$ to obtain
$\left(\pi(D_{p})+\pi(D_{4p})\right)-2nc_{\infty}^{+}$
where $c_{\infty}^{+}$ is the unique cusp of $X_{0}^{+}(p)$. Applying Proposition 12 we get:
$h\left(-4p\right)+h\left(-p\right)\geq\frac{p}{96}$.
When $p\equiv 7\mod{8}$, $h\left(-4p\right)=h\left(-p\right)$ and we find that for $p\geq 749,504$:
$2h\left(-p\right)\leq\frac{2}{3}\sqrt{p}\log p<\frac{p}{96}$.
Computationally, we reduce the above bound to $104711$ and note that there are $566$ primes satisfying the stated congruence conditions.
When $p\equiv 3\mod{8}$, $h\left(-4p\right)=3h\left(-p\right)$, and repeating the above we get $p\geq 3754815$, which we reduce to $49451$, and $266$ primes to check. The list of exceptions includes the primes $\{11,23,47,59,71\}$, for which $X_{0}(p)$ is hyperelliptic. For all non-hyperelliptic exceptions we find a supersingular $j$-invariant defined over $\mathbb{F}_{p^{2}}\setminus\mathbb{F}_{p}$ and argue as in Proposition 20. These cases are checked in the file primes11AL.m, using that the $j$-invariants of points in $D_{p}$ and $D_{4p}$ reduce to the $\mathbb{F}_{p}$-rational supersingular $j$-invariants modulo $p$.
5. Proof of Theorem 2
In this section we prove that the Ceresa cycle is non-trivial for all but finitely many $X_{0}(N)$, where $N$ is any positive integer. Our proof uses the following elementary result.
Proposition 24.
Let $C$ and $D$ be algebraic curves of genus at least $3$, such that $C$ covers $D$ and $\operatorname{Cer}(D,f)$ is non-zero in $CH_{1}(J)$ for all choices of $f$. Then $\operatorname{Cer}(C,e)$ is also non-zero for any choice of basepoint divisor $e$.
Proof.
Let $\pi:C\to D$ denote the covering map. As $\pi_{*}[\iota_{e}(C)]=\deg(\pi)[\iota_{\pi(e)}(D)]$, we have $\pi_{*}\operatorname{Cer}(C,e)=\deg(\pi)\operatorname{Cer}(D,\pi(e))$. By assumption, $\operatorname{Cer}(D,f)$ is non-torsion for all base-points, $f$, and so the pushforward of $\operatorname{Cer}(C,e)$ is of infinite order, and in particular, $\operatorname{Cer}(C,e)$ is also of infinite order. ∎
For any prime $p$ dividing $N$, the natural quotient map induces a covering of curves
$X_{0}(N)\longrightarrow X_{0}\left(p\right)$
and thus from Theorem 1, it follows that if $N$ is divisible by a prime $p$ and
$p\notin\{2,3,5,7,11,13,17,19,23,29,31,37,41,47,59,71\}$
that is, $X_{0}(p)$ has genus at least $3$ and it is not hyperelliptic (see [31]), then $\operatorname{Cer}(N)$ is of infinite order. It remains to show that for any $p$ in the above list, $\operatorname{Cer}(p^{n})$ is non-torsion for some $n\geq 2$. For most primes we take $n=2$. When $p=2,3,5,7$, we take larger powers to ensure that the Jacobian is not hyperelliptic and usually also of positive rank.
The case for $p=2$ is well known.
Theorem 25.
The cycle $\operatorname{Cer}(64)$ is non-zero.
Proof.
It is a well-known fact that the modular curve $X_{0}(64)$ is the Fermat quartic. Non-vanishing is verified directly in [16]. Moreover, in [21] Harris famously proved that the Ceresa cycle of the Fermat quartic is not algebraically equivalent to zero. ∎
5.1. Heegner Divisors
For some $N\equiv 11\mod{12}$, certain shadows are supported on Heegner points and cusps only, in which case we deduce that the (scaled) shadow point is non-torsion using the classical theory of Heegner divisors.
Let $P\in X_{0}(N)$ be a Heegner point (that is, a CM point which satisfies the Heegner hypothesis). The Heegner divisor defined by $P$ is the degree 00 divisor
$D_{P}=\displaystyle\sum_{\sigma}P^{\sigma}-Mc_{\infty}$
where the sum is taken over the Galois conjugates of $P$, $c_{\infty}$ is the $\infty$ cusp of $X_{0}(N)$ and $M$ is the degree of the field over which $P$ is defined. The following is a simple consequence of results proved in [19](or more generally [20]).
Theorem 26 (Gross).
Let $P$ and $D_{P}$ be as above. Then $[D_{P}]$ is a point of infinite order in $J_{0}(N)(\mathbb{Q})$.
For $p\in\{23,47,59,71\}$, we take $N=p^{2}$ and note that $X_{0}(N)$ has $p+1$ cusps, $c_{\infty},c_{0},c_{1},\ldots,c_{p-1}$, where $c_{\infty}$ and $c_{0}$ are defined over $\mathbb{Q}$ and $c_{1},\ldots,c_{p-1}$ form a single Galois orbit. We write $D$ for the divisor class corresponding to the sum of the cusps. We compute the following shadows:
  • $N=23^{2}:6\operatorname{Sh}(T_{2})=816D_{7}-68D$;
  • $N=47^{2}:6\operatorname{Sh}(T_{3})=3936D_{11}-164D$;
  • $N=59^{2}:6\operatorname{Sh}(T_{5})=6360D_{11}-212D$;
  • $N=71^{2}:6\operatorname{Sh}(T_{2})=9360D_{7}-260D$.
The scaled shadow points differ from a Heegner divisor by a divisor supported on cusps only, and as such divisors are torsion points in $J_{0}(N)$ by the theorems of Manin [28] and Drinfeld [15], we conclude that the scaled shadow points are necessarily non-torsion.
5.2. Explicit Calculations
For curves of low genus, we can use Magma to efficiently compute divisors on the curve and explicitly test whether our chosen shadow point is torsion.
For $N\in\{11^{2}$, $5^{3}$, $13^{2}\}$, we work with $D=\operatorname{Sh}(w_{N})$. We begin by computing a canonical model for the curve using a basis of the space of weight $2$ cusps forms for $\Gamma_{0}(N)$, see [18]. In particular, we choose a basis whose elements are invariant or anti-invariant under the action of $w_{N}$ and from this we are able easily construct $D$ explicitly. We show that modulo different primes, the divisor has different orders, and hence by the injectivity of torsion [22, Appendix], we conclude that the shadow is necessarily of infinite order. See the online repository for these explicit calculations.
Moreover, we also verify that $X_{0}(74)$ has non-trivial Ceresa cycle by the same methods. This, along with Theorem 1 and the proof of Theorem 2 recovers the main result of [23].
5.3. Period Lattice Computations
For $N\in\{3^{5},7^{3},17^{2},19^{2},29^{2},31^{2},37^{2},41^{2}\}$, the genus of $X_{0}(N)$ is large and computing explicit divisors in Magma is extremely inefficient. Instead, we use the analytic structure of the curve and its Jacobian, and project our shadow point into a (complex) subvariety of dimension $1$, $2$ or $3$. We give a short overview of the analytic definition of the Jacobian $J_{0}(N)$ and its modular subvarieties, see [13, Chapter 6] or [8, Chapter 2] for comprehensive descriptions.
Let $f_{1},\ldots,f_{g}$ be a basis for $S_{2}(N)$, the space of weight 2 cusp forms for $\Gamma_{0}(N)$, and take $\gamma_{1},\ldots,\gamma_{2g}$ to be a basis for the integral singular homology $H_{1}(X_{0}(N),\mathbb{Z})$. The period matrix defined with respect to these bases is the $2g\times g$ complex matrix
$\Omega=(\omega_{jk})=(\langle\gamma_{j},f_{k}\rangle)$
[…]
Any cusp form $f\in S_{2}(N)$ has a Fourier expansion of the form
$f(\tau)=\displaystyle\sum_{n=1}^{\infty}a_{n}e^{2\pi i\tau}$
where the coefficients $a_{n}\in\mathbb{C}$ are in fact algebraic integers. Let $g_{1},\ldots,g_{d}$ be the set of Galois conjugates of $f$. The period lattice of $f$ is the $\mathbb{Z}$-span of the vectors
$(\langle\gamma_{i},g_{1}\rangle,\ldots,\langle\gamma_{i},g_{d}\rangle)$
where $1\leq i\leq 2g$. This is a lattice of rank $2d$ in $\mathbb{C}^{d}$, and the quotient $A_{f}=\mathbb{C}^{d}/\Lambda$ is an abelian variety defined over $\mathbb{Q}$. Moreover, there is a natural quotient map
$\varphi_{f}:J_{0}(N)\longrightarrow A_{f}$
which is also defined over $\mathbb{Q}$. Notably, when $d=1$, $A_{f}$ is simply the elliptic curve associated to $f$ and $\varphi_{f}$ is the modular parametrisation.
For $N\in\{3^{5},17^{2},19^{2},29^{2},37^{2},41^{2}\}$, we note that $S_{2}(N)$ contains a positive-rank newform $f$: of dimension $1$ for $N\in\{3^{5},17^{2},19^{2},37^{2}\}$; of dimension $2$ for $N\in\{29^{2},31^{2}41^{2}\}$; and of dimension $3$ for $N=7^{3}$. Let $D$ be a chosen scaled shadow which we assume to be torsion in $J_{0}(N)$ (we use the shadow corresponding to $T_{2}$, except for $7^{3}$ where we use $T_{3}$). Then, for any upper bound $B(f)$ of $|A_{f}(\mathbb{Q})_{{\operatorname{tors}}}|$, $\varphi_{f}([B(f)\cdot D])=0$ and hence $B(f)\cdot D\in\Lambda_{f}$. Showing that this does not occur reduces to a finite computation as we now explain.
Firstly, bounding $|A_{f}(\mathbb{Q})_{{\operatorname{tors}}}|$ is elementary, as for any odd prime $l\nmid N$, reduction modulo $l$ induces an injection
$A_{f}(\mathbb{Q})_{\text{tors}}\longrightarrow A_{f}(\mathbb{F}_{l})$
see [22, Appendix]. We may therefore take $B(N)$ to be the lowest common multiple of $|A_{f}(\mathbb{F}_{l})|$, as $l$ ranges over a finite set of odd primes coprime to $N$. Moreover, the cardinality of $A_{f}(\mathbb{F}_{l})$ can be determined directly from the coefficients of $f$ as detailed in [1, Section 3.5].
Let $D$ be $D^{\prime}-C$, where $C$ is supported on the cusps and $D^{\prime}$ is supported away from the cusps. The divisor $\tilde{D}=D^{\prime}-\deg(D^{\prime})c_{\infty}$, where $c_{\infty}\in X_{0}(N)$ is the infinity cusp, differs from $D$ by a divisor which is supported on cusps. By the theorems of Manin [28] and Drinfeld [15], the difference between $D$ and $\tilde{D}$ is torsion, and so $D$ is of infinite order if and only if $\tilde{D}$ is. We therefore prove $\tilde{D}$ is non-torsion, equivalently, $B(f)\tilde{D}\notin\Lambda_{f}$. Note that the CM points $\tau_{i}\in\mathbb{H}$ corresponding to the CM points $P_{i}$ in the support of the $D^{\prime}$ can be explicitly determined in our examples.
Let $D^{\prime}=\sum_{i}a_{i}P_{i}$ and construct an integral basis $I$ of the subspace of $S_{2}(N)$ spanned by $f$ and its conjugates. In our calculations, $I$ has at most $3$ elements. Any element $h\in I$ has a Fourier expansion of the form
$h(\tau)=\displaystyle\sum_{n\geq 0}b_{n}^{h}e^{2\pi i\tau n}$
with $b_{i}^{h}\in\mathbb{Z}$; and we use this to verify that
$\left(\displaystyle\sum_{i=1}a_{i}\int_{\infty}^{\tau_{i}}2\pi ih(z)dz:h\in I\right)\not\in\Lambda_{f}$.
In particular, from Deligne’s proof of the Weil conjectures [11], we deduce that $|Cb_{n}^{h}|\leq d(n)n^{1/2}$, where $d(n)$ denotes the number of positive divisors of $n$ and $C$ is an explicit constant depending on how $h$ is expressed in terms of the basis of normalised eigenforms. We use this to bound the absolute value of the integral of the truncated series
$h_{m}\left(\tau\right)=\displaystyle\sum_{n=1}^{m}b_{n}^{h}e^{2\pi i\tau n}$
by the geometric series $\sum_{n=1}^{m}2e^{2n\pi i\tau}$. Summing this geometric series we obtain an estimate for the error, and so we can deduce a choice of $m$ which guarantees convergence to within a chosen tolerance. This allows us to approximate the value of the integral. We compute the period lattice associated to $f$ to within the same tolerance using the Magma command Periods(), see [8] for details of this algorithm. The final step is a simple linear algebra computation which verifies that the point does not lie in the lattice.
These calculations can often be time consuming due to the need to compute many terms of the Fourier expansions, and Hecke operator calculations become increasingly expensive. The worst case was $41^{2}$, which took over 12 hours due to needing the first 6600 terms of the expansion.